For three events , and , if (exactly one of or occurs) (exactly one of or occurs) (exactly one of or occurs) and (all the three events occur simultaneously) , then the probability that atleast one of the events occurs, is
- A.
- B.
- C.
- D.
Show answer and solution
Correct option: B
Solution
The key identity: P(exactly one of A or B) = P(A) + P(B) - 2P(A∩B) = 1/4.
Similarly for the other pairs. Adding all three equations:
2[P(A) + P(B) + P(C)] - 2[P(A∩B) + P(B∩C) + P(C∩A)] = 3/4
So P(A) + P(B) + P(C) - P(A∩B) - P(B∩C) - P(C∩A) = 3/8.
Now use inclusion-exclusion for P(at least one):
P(A∪B∪C) = P(A)+P(B)+P(C) - P(A∩B) - P(B∩C) - P(C∩A) + P(A∩B∩C)
= 3/8 + 1/16
= 6/16 + 1/16 = 7/16.
Similarly for the other pairs. Adding all three equations:
2[P(A) + P(B) + P(C)] - 2[P(A∩B) + P(B∩C) + P(C∩A)] = 3/4
So P(A) + P(B) + P(C) - P(A∩B) - P(B∩C) - P(C∩A) = 3/8.
Now use inclusion-exclusion for P(at least one):
P(A∪B∪C) = P(A)+P(B)+P(C) - P(A∩B) - P(B∩C) - P(C∩A) + P(A∩B∩C)
= 3/8 + 1/16
= 6/16 + 1/16 = 7/16.
[JEE Main 2017]